OpenStudy (dako87):

What is the slope of the line passing through the points (-4, 3) and (5, -3)?

OpenStudy (greatlife44):

\[m~slope = \frac{ y_{2}-y_{1} }{ x_{2}-x_{2} }\]

RhondaSommer (rhondasommer):



OpenStudy (dako87):

-6/9 is what i got

OpenStudy (greatlife44):

so you would first subtract the y and x coordinates like this y/x = (-3-3)/(5-(-4)) = ?

OpenStudy (mathstudent55):